If now Lina is three times as old as Nick, and in 6 years she will be twice as old as he, how old are they now?

Answer by Alan3354(69443)

If now Lina is three times as old as Nick, and in 6 years she will be twice as old as he, how old are they now?
------------
18 & 6, by thinking about it for 2 seconds.
====================
The hard way:
L = 3N --- ages now
---
L + 6 = 2*(N + 6)
3N + 6 = 2N + 12
N = 6 --- Nick's age
L = 18 -- Lina's age